Operations Report – April 23rd

Crew 193 Operations Report 23Apr2018

SOL: 1

Name of person filing report: David Attig

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): 14

Solar— SOC 92 % (Before generator is run at night)

Diesel – 75 %

Propane – 60 psi.

ATV Fuel (Ethanol free gasoline) – 1 gallons.

Water (trailer) – 0 gallons.

Water (storage) – 550 gallons

Water (static) – 470 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 133949.8

Toilet tank emptied: no
Date last emptied: April 21

ATV’s (Honda, 300, 350.1, 350.2, 350.3):
Honda Used?: Yes
300 Used? No
350.1 Used? Yes
350.2 Used? Yes
350.3 Used? Yes

ATV Oil Checked?: No

ATV Oil Added?: No

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0.3

Notes on ATVs: None

Deimos rover used: No
Hours: 123.5
Beginning charge: 100%
Ending charge: 100 %
Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unknown
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: No
Hours: 33.7
Beginning charge: 100 %
Ending charge: 100 %
Currently charging: No

Opportunity rover used: Out Of Service
Hours: 31
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: No
Hours: 23
Beginning charge: 100 %
Ending charge: 100 %
Currently charging: No

Hab Car used?: No
Where was Hab Car used?: N/A
Why was Hab Car used?: N/A

General notes and comments: This morning we had an issue with the readings on the solar SoC before we turned off the generator. Systems seem to be nominal now.

Summary of internet: Nominal. This morning the internet was very slow, while we still had 49% of our

Summary of suits and radios: All one piece suits are working (and were used today). Radios are nominal.

Summary of Hab operations: See Sol Summary report.

Summary of GreenHab operations: See GreenHab report

Summary of ScienceDome operations: 001 constructed dishwasher experiment. While checking SoC, I noticed one of the nuts had fallen from lock. It has been replaced, but is not fully tightened as that would interfere with Lock operations. We will keep an eye on it, suggest a longer bolt with two nuts jammed together in future upgrades.

Summary of RAM operations: Sign marking area of our settlement was in RAM. Old sign was repaired and moved outside. New sign material is in RAM awaiting construction. Door was left open, engineering EVA was performed to close it.

Summary of health and safety issues: None to report.

Questions, concerns and requests to Mission Support: Nothing yet.

Copyright © The Mars Society. All rights reserved. | Main Site